# Cybersecurity of Quantum Key Distribution Implementations

### v2

@@ C.3 Attack Constraints
Appendix D Additional Examples for Reversed-Space Attacks
In this appendix, we give additional examples for the application of the ideas from Section 5 on different setups, especially on the computation of the classical interpretation $J$ sets and the process of finding an oblivious attack.
D.1 Examples of Interpretation Sets in Imperfect Bob Setups
Recall the example of measurement imperfections in Section 3.2.1 where Bob receives a pulse of photons with uniform polarization, written as the Fock space state $|{m_{1},m_{0}}\rangle^{\!\text{\tiny F}}$ with $m_{1},m_{0}\geq 0$. Let us assume that Bob’s detector cannot distinguish one photon from two or more photons, i.e., any state $|{0,k}\rangle^{\!\text{\tiny F}}$ with $k\geq 1$ causes the “0” detector to click, while any state $|{k,0}\rangle^{\!\text{\tiny F}}$ with $k\geq 1$ causes the “1” detector to click. To make the analysis easier, we limit the discussion below to pulses with at most two photons.
In a BB84 implementation using this setting, Bob chooses to either measure using the computational basis (rectilinear polarization) or the Hadamard basis (diagonal polarization). We denote these settings by $s=1$ and $s=2$, respectively. To measure in the computational basis, Bob simply lets the state go into his detector; to measure in the Hadamard basis, Bob applies the Hadamard gate on the pulse and then measures the outcome.
The Hadamard gate evolves the possible 6 states (i.e., pulses with up to 2 photons) as
$(1)\ |{0,0}\rangle^{\!\text{\tiny F}}$ $\to|{0,0}\rangle^{\!\text{\tiny F}}$ $(4)\ |{1,1}\rangle^{\!\text{\tiny F}}$ $\to(|{0,2}\rangle^{\!\text{\tiny F}}-|{2,0}\rangle^{\!\text{\tiny F}})/\sqrt{2}$
$(2)\ |{0,1}\rangle^{\!\text{\tiny F}}$ $\to(|{0,1}\rangle^{\!\text{\tiny F}}+|{1,0}\rangle^{\!\text{\tiny F}})/\sqrt{2}$ $(5)\ |{0,2}\rangle^{\!\text{\tiny F}}$ $\to(|{0,2}\rangle^{\!\text{\tiny F}}+\sqrt{2}|{1,1}\rangle^{\!\text{\tiny F}}+|{2,0}\rangle^{\!\text{\tiny F}})/2$
$(3)\ |{1,0}\rangle^{\!\text{\tiny F}}$ $\to(|{0,1}\rangle^{\!\text{\tiny F}}-|{1,0}\rangle^{\!\text{\tiny F}})/\sqrt{2}$ $(6)\ |{2,0}\rangle^{\!\text{\tiny F}}$ $\to(|{0,2}\rangle^{\!\text{\tiny F}}-\sqrt{2}|{1,1}\rangle^{\!\text{\tiny F}}+|{2,0}\rangle^{\!\text{\tiny F}})/2.$
Using the same numbering for basis state, for the $s=1$ setting we have $\beta^{s=1}_{[1,6]\times[1,6]}=I$, and for the $s=2$ setting,
$\beta^{s=2}_{[1,6]\times[1,6]}=\begin{pmatrix}1&0&0&0&0&0\\ 0&1/\sqrt{2}&1/\sqrt{2}&0&0&0\\ 0&1/\sqrt{2}&-1/\sqrt{2}&0&0&0\\ 0&0&0&0&1/\sqrt{2}&-1/\sqrt{2}\\ 0&0&0&1/\sqrt{2}&1/{2}&1/{2}\\ 0&0&0&-1/\sqrt{2}&1/{2}&1/{2}\\ \end{pmatrix}$.
We can assume that Bob considers the case where both detectors click as an invalid state, yet if no detector clicks, Bob considers this case as a valid loss. Then, using the notations presented in Section 5.2, $J_{\operatorname{loss}}=\{|{0,0}\rangle^{\!\text{\tiny F}}\}$ and $J_{\operatorname{invalid}}=\{|{1,1}\rangle^{\!\text{\tiny F}}\}$. The set $J_{\operatorname{error}}$ depends on the qubit sent by Alice. If Alice sends a 0 (either using the computational or the Hadamard basis) then a click of the “1”-detector is an error, and since the detector does not distinguish two photons from one we have $J_{{\operatorname{error}}}=J_{1}=\{|{1,0}\rangle^{\!\text{\tiny F}},|{2,0}\rangle^{\!\text{\tiny F}}\}$. Similarly, if Alice sends a 1 then $J_{{\operatorname{error}}}=J_{0}=\{|{0,1}\rangle^{\!\text{\tiny F}},|{0,2}\rangle^{\!\text{\tiny F}}\}$.
Bob adds no ancilla in this case, and his entire measured space arrives from the channel. Thus, from the reversed space analysis we get $\mathcal{H}^{\mathrm{B}}=\mathcal{H}^{\mathrm{P}}$. Note that the reversed space in this case has a larger dimension than the theoretical qubit space sent by Alice: it also includes pulses with zero photons and pulses with two photons.
If we do not limit ourselves to attacks in which Eve sends at most two photons, we can recover a well known attack: for this, we need to add the assumption that if both Bob’s detectors click, he considers this invalid outcome as a loss and ignores it. In this case, Eve can perform a measure-and-resend attack in one of the bases, such that when she resends, the state is the same as the one she measured, but with a much larger number of photons (this is an example of the “faked states” [19, 20] attacks described in Section 3.2.3). For example, if Eve measures $|{0,1}\rangle^{\!\text{\tiny F}}$, she sends $|{0,m}\rangle^{\!\text{\tiny F}}$ with $m\gg 1$. Given that Bob’s detector do not distinguish one photon from many, if Bob uses the same basis as Eve, he gets a legitimate result, and he never gets errors; and if he measures in the other basis, many photons enter both detectors (with high probability), both detectors click, and Bob ignores this outcome. This attack allows Eve to get full information without being noticed [55, 60].
D.2 Insecurity of Interferometric BB84: When Bob Does Not Measure All Modes
Consider the case in which Bob wants to open his detector only during a single detection slot. This may be done in order to achieve a higher bit-rate, or may be forced due to technological (or financial) limitations that restrict Bob from opening the detectors to more than a single detection window per pulse. [^9] This issue is usually relevant in telecommunication wave length (IR spectrum) technology. A possible scheme for such a limited Bob is to measure only a single time-bin in each detector. For instance, to perform a measurement in the computational basis, assume Bob measures only $\{|{d_{0}}\rangle,|{s_{2}}\rangle\}$, i.e., opening the $d$-arm detector at time $t_{0}$ (to measure $|{0}\rangle$) and the $s$ arm detector at time $t_{2}$ (to measure $|{1}\rangle$). For measuring in the Hadamard basis, Bob opens both his detectors at time $t_{1}$: that is, he measures $|{d_{1}}\rangle,|{s_{1}}\rangle$ to indicate $|{+}\rangle$ and $|{-}\rangle$, respectively. This (practical) weakening of Bob leads to a stronger attack in which the detection efficiency in each basis is reduced by half; yet, in contrast to the attack described in Section 5.3 and Appendix B, it is not the case where one basis will always yield a loss.
Using the tools presented in Section 5.2, we find $\mathbf{U}_{\mathrm{zero}}$ (defined in Definition 7 as the set of attacks not causing any error) and discover that it includes attacks capable of revealing the information in its entirety to Eve. It should be noted that, as in Appendix B, the attacks are individual-particle attacks in which Eve uses only single-photon pulses; limiting Eve to single-photon pulses simplifies the analysis and is sufficient for proving insecurity.
Note that a basis of $\mathcal{H}^{\mathrm{B}}$ in this restricted case is defined as $B=\{|{V}\rangle,|{d_{0}}\rangle,|{d_{1}}\rangle,|{s_{1}}\rangle,|{s_{2}}\rangle\}$. The reversed space $\mathcal{H}^{\mathrm{P}}$ in this case (following Eq. (27) and Section B.2) is the span of $\{|{V}\rangle$, $|{t^{\prime}_{-1}}\rangle$, $|{t^{\prime}_{0}}\rangle$, $|{t^{\prime}_{1}}\rangle$, $|{t^{\prime}_{2}}\rangle\}$.
Bob interprets his measurement outcome in the following manner: for the Hadamard basis, $J_{0}=\{|{d_{1}}\rangle\}$; $J_{1}=\{|{s_{1}}\rangle\}$; $J_{\operatorname{loss}}=B\setminus(J_{0}\cup J_{1})$; and $J_{\operatorname{invalid}}=\emptyset$. For the computational basis, Bob uses the following interpretation: $J_{0}=\{|{d_{0}}\rangle\}$; $J_{1}=\{|{s_{2}}\rangle\}$; $J_{\operatorname{loss}}=B\setminus(J_{0}\cup J_{1})$; the set $J_{\operatorname{invalid}}$ is empty again due to the assumption of using only single-photon pulses.
As done in Appendix B, we use Observation 1 to define the set of Eve’s attacks that cause no errors. The requirements for the Hadamard basis are the same as in Section B.3 and are given by Eqs. (28)–(29). The requirements for the computational basis are
$\epsilon_{t^{\prime}_{0},t^{\prime}_{1}}|{E_{t^{\prime}_{0},t^{\prime}_{1}}}\rangle-\epsilon_{t^{\prime}_{0},t^{\prime}_{2}}|{E_{t^{\prime}_{0},t^{\prime}_{2}}}\rangle=0$ when Alice sends $|{0}\rangle$, and (57)
$i\epsilon_{t^{\prime}_{1},t^{\prime}_{-1}}|{E_{t^{\prime}_{1},t^{\prime}_{-1}}}\rangle+i\epsilon_{t^{\prime}_{1},t^{\prime}_{0}}|{E_{t^{\prime}_{1},t^{\prime}_{0}}}\rangle=0$ when Alice sends $|{1}\rangle$. (58)
Thus, the family of Eve’s attacks that cause no errors is of the form (omitting the vacuum state)
$|{0}\rangle_{\mathrm{E}}|{0}\rangle_{\mathrm{A}}$ $\stackrel{{\scriptstyle U_{\mathrm{E}}}}{{\longrightarrow}}\phantom{-}p_{1}|{E_{1}}\rangle_{\mathrm{E}}|{t^{\prime}_{-1}}\rangle_{\mathrm{P}}+p_{2}|{E_{2}}\rangle_{\mathrm{E}}|{t^{\prime}_{0}}\rangle_{\mathrm{P}}+p_{3}|{E_{3}}\rangle_{\mathrm{E}}|{t^{\prime}_{1}}\rangle_{\mathrm{P}}+p_{3}|{E_{3}}\rangle_{\mathrm{E}}|{t^{\prime}_{2}}\rangle_{\mathrm{P}}$
$|{0}\rangle_{\mathrm{E}}|{1}\rangle_{\mathrm{A}}$ $\stackrel{{\scriptstyle U_{\mathrm{E}}}}{{\longrightarrow}}-p_{3}|{E_{3}}\rangle_{\mathrm{E}}|{t^{\prime}_{-1}}\rangle_{\mathrm{P}}+p_{3}|{E_{3}}\rangle_{\mathrm{E}}|{t^{\prime}_{0}}\rangle_{\mathrm{P}}+p_{2}|{E_{2}}\rangle_{\mathrm{E}}|{t^{\prime}_{1}}\rangle_{\mathrm{P}}+p_{4}|{E_{4}}\rangle_{\mathrm{E}}|{t^{\prime}_{2}}\rangle_{\mathrm{P}},$
with the normalization conditions $|p_{1}|^{2}+|p_{2}|^{2}+2|p_{3}|^{2}=|p_{4}|^{2}+|p_{2}|^{2}+2|p_{3}|^{2}=1$.
It follows that Alice’s qubit evolves as
$|{0}\rangle_{\mathrm{A}}\to\frac{p_{1}|{E_{1}}\rangle}{2}|{s_{-1}}\rangle+\frac{p_{2}|{E_{2}}\rangle-p_{1}|{E_{1}}\rangle}{2}|{s_{0}}\rangle+\frac{p_{3}|{E_{3}}\rangle-p_{2}|{E_{2}}\rangle}{2}|{s_{1}}\rangle+\frac{p_{3}|{E_{3}}\rangle}{2}|{s_{3}}\rangle+i\frac{p_{1}|{E_{1}}\rangle}{2}|{d_{-1}}\rangle\\ +i\frac{p_{2}|{E_{2}}\rangle+p_{1}|{E_{1}}\rangle}{2}|{d_{0}}\rangle+i\frac{p_{3}|{E_{3}}\rangle+p_{2}|{E_{2}}\rangle}{2}|{d_{1}}\rangle+2i\frac{p_{3}|{E_{3}}\rangle}{2}|{d_{2}}\rangle+i\frac{p_{3}|{E_{3}}\rangle}{2}|{d_{3}}\rangle$
$|{1}\rangle_{\mathrm{A}}\to\frac{-p_{3}|{E_{3}}\rangle}{2}|{s_{-1}}\rangle+2\frac{p_{3}|{E_{3}}\rangle}{2}|{s_{0}}\rangle+\frac{p_{2}|{E_{2}}\rangle-p_{3}|{E_{3}}\rangle}{2}|{s_{1}}\rangle+\frac{p_{4}|{E_{4}}\rangle-p_{2}|{E_{2}}\rangle}{2}|{s_{2}}\rangle-\frac{p_{4}|{E_{4}}\rangle}{2}|{s_{3}}\rangle\\ +i\frac{p_{3}|{E_{3}}\rangle}{2}|{d_{-1}}\rangle+i\frac{p_{2}|{E_{2}}\rangle+p_{3}|{E_{3}}\rangle}{2}|{d_{1}}\rangle+i\frac{p_{2}|{E_{2}}\rangle+p_{4}|{E_{4}}\rangle}{2}|{d_{2}}\rangle+i\frac{p_{4}|{E_{4}}\rangle}{2}|{d_{3}}\rangle,$
and never causes an error in the computational basis. For the Hadamard basis, $|{+}\rangle_{\mathrm{A}}$ and $|{-}\rangle_{\mathrm{A}}$ follow from the above by linearity. For concreteness, we only note that the projection on the states Bob measure behaves as required, i.e.,
$|{+}\rangle_{\mathrm{A}}$ $\to 0\cdot|{s_{1}}\rangle+\frac{i(p_{3}|{E_{3}}\rangle+p_{2}|{E_{2}}\rangle)}{\sqrt{2}}|{d_{1}}\rangle+\ldots$
$|{-}\rangle_{\mathrm{A}}$ $\to\frac{p_{3}|{E_{3}}\rangle-p_{2}|{E_{2}}\rangle}{\sqrt{2}}|{s_{1}}\rangle+0\cdot|{d_{1}}\rangle+\ldots$
Let us define the variable $r$ as the bit value measured by Bob (restricted to the case where Bob and Alice use the same basis, and no loss has occurred). When Bob’s outcome is $r$, Eve holds a state as described in Table 1.
Alice’s State Eve’s State / $r=0$ Eve’s State / $r=1$
$|{{0}}\rangle_{\mathrm{A}}$ $p_{1}|{E_{1}}\rangle_{\mathrm{E}}+p_{2}|{E_{2}}\rangle_{\mathrm{E}}$
$|{{1}}\rangle_{\mathrm{A}}$   $p_{4}|{E_{4}}\rangle_{\mathrm{E}}-p_{2}|{E_{2}}\rangle_{\mathrm{E}}$
$|{{+}}\rangle_{\mathrm{A}}$ $p_{3}|{E_{3}}\rangle_{\mathrm{E}}+p_{2}|{E_{2}}\rangle_{\mathrm{E}}$
$|{{-}}\rangle_{\mathrm{A}}$   $p_{3}|{E_{3}}\rangle_{\mathrm{E}}-p_{2}|{E_{2}}\rangle_{\mathrm{E}}$
Table 1: The non-normalized states Eve holds as a function of the state sent by Alice, conditioned on Bob measuring the bit value $r$ (and not a loss)
Eve can acquire full information about the original state, for instance, by setting $p_{i}=1/2$, $|{E_{1}}\rangle=|{E_{4}}\rangle$, and letting $|{E_{i}}\rangle_{\mathrm{E}}$ be orthogonal for $i=1,2,3$. Once Alice reveals the basis used, Eve measures her state and reveals Alice’s bit with certainty. However, this reduces Bob’s detection efficiency: when no attack is present, Bob measures a valid value with probability $1/2$ for the computational basis, and probability $1/2$ for the Hadamard basis. With the above attack, Bob’s detection efficiency decreases to $1/8$ in the computational basis and $1/4$ in the Hadamard basis. By setting the parameters $\{p_{i}\}_{i=1,2,3,4}$, Eve can consider a tradeoff between the amount of information she acquires and the loss rate she induces on Bob.
Appendix E Formal Definition of Quantum Side-Channel Attacks
This is the formal formulation of Definitions 8 and 9, presented in Section 6.
  • Let $\mathcal{H}^{\mathrm{A}}_{\mathrm{full}}$ be the Hilbert space of all states that can arrive from Alice’s channel, and let $\mathcal{H}^{\mathrm{B}}_{\mathrm{full}}$ be the Hilbert space of all states that can arrive at Bob’s channel.
  • Let $\mathcal{H}^{\mathrm{A}}\subseteq\mathcal{H}^{\mathrm{A}}_{\mathrm{full}}$ be the span of Alice’s sent states, and let $\mathcal{H}^{\mathrm{B}}\subseteq\mathcal{H}^{\mathrm{B}}_{\mathrm{full}}$ be the space of states that affect Bob’s measurement.
  • Let $\mathcal{H}^{\mathrm{env}}$ be the Hilbert space representing the publicly available environment that is not part of the previously defined spaces.
  • Let Eve’s attack be defined as the following unitary transformation:
    $U_{\mathrm{E}}:\mathcal{H}^{\mathrm{A}}_{\mathrm{full}}\otimes\mathcal{H}^{\mathrm{anc}}_{\mathrm{in}}\otimes\mathcal{H}^{\mathrm{env}}\;\to\;\mathcal{H}^{\mathrm{B}}_{\mathrm{full}}\otimes\mathcal{H}^{\mathrm{anc}}_{\mathrm{out}}\otimes\mathcal{H}^{\mathrm{env}},$
    where $\mathcal{H}^{\mathrm{anc}}_{\mathrm{in}},\mathcal{H}^{\mathrm{anc}}_{\mathrm{out}}$ are Eve’s private ancilla spaces.
Definition 10.
Eve’s attack is a Quantum Side-Channel Attack if either of the following two conditions is true:
  1. 1.
    Eve’s attack depends non-trivially on $\mathcal{H}^{\mathrm{env}}$;
  2. 2.
    the output of Eve’s attack non-trivially affects either $\mathcal{H}^{\mathrm{B}}_{\mathrm{full}}\setminus\mathcal{H}^{\mathrm{B}}$ or $\mathcal{H}^{\mathrm{env}}$.
Definition 11.
Eve’s attack is a Quantum State-Channel Attack if the following two conditions are both true:
  1. 1.
    Eve’s attack does not depend non-trivially on $\mathcal{H}^{\mathrm{env}}$;
  2. 2.
    the output of Eve’s attack non-trivially affects only $\mathcal{H}^{\mathrm{B}}$ and $\mathcal{H}^{\mathrm{anc}}_{\mathrm{out}}$.
Appendix F Computation of Bright Illumination Attack Using Reversed-Space
In this appendix, we show the full computation of the Reversed-Space Attack shown in Section 7.4, beginning from the behavior of the implementation that was discovered through Quantum Fuzzing in Section 7.3 and resulting in the original attack described in Section 7.1. Our modeling of the device will rely only on the information learned on the device from fuzzing, and on the fact that the device employs a single unitary transformation independently of the basis it wishes to measure.
@@ Appendix F Computation of Bright Illumination Attack Using Reversed-Space
The vulnerabilities shown in Section 7.3 show that Bob’s device has an unintended configuration that can be triggered by Eve (the linear mode of the APDs), and that in this configuration, his device suffers from a collection of Measurement Space Vulnerabilities and Interpretation Vulnerabilities (defined in Section 3.2.1). We will now formally describe these vulnerabilities.
@@
Note that, according to the definition of high-photon-number states with diagonal polarization, the overlap between “bright Hadamard” states and “bright computational” states is given, without loss of generality, by:
${\langle{+}|}^{\mathrm{bright}}|{0}\rangle^{\mathrm{bright}}=\left(\frac{1}{\sqrt{2^{k}}}\sum_{l=0}^{k}\sqrt{\binom{k}{l}}{\langle{l,k-l}|}_{\mathrm{H},\mathrm{V}}\right)|{0,k}\rangle_{\mathrm{H},\mathrm{V}}=\frac{1}{\sqrt{2^{k}}}=O\left(2^{-\frac{k}{2}}\right).$ (60)
@@
$\mathcal{H}^{\mathrm{B}}=\operatorname{span}\left\{|{\psi_{0}}\rangle_{\mathrm{B}},|{\psi_{1}}\rangle_{\mathrm{B}},|{\psi_{2}}\rangle_{\mathrm{B}},|{\psi_{3}}\rangle_{\mathrm{B}}\right\}.$ (61)
@@
    $J_{0}=\{|{\psi_{0}}\rangle_{\mathrm{B}}\},\quad J_{1}=\{|{\psi_{1}}\rangle_{\mathrm{B}}\}.$ (62)
@@
    $J_{0}=\{|{\psi_{2}}\rangle_{\mathrm{B}}\},\quad J_{1}=\{|{\psi_{3}}\rangle_{\mathrm{B}}\}.$ (63)
@@
$\mathcal{H}^{\mathrm{P}}=\operatorname{span}\left\{U_{\mathrm{B}}^{\dagger}|{\psi_{i}}\rangle_{\mathrm{B}}\;\big{|}\;i=0,1,2,3\right\}=\operatorname{span}\left\{|{0}\rangle^{\mathrm{bright}},|{1}\rangle^{\mathrm{bright}},|{+}\rangle^{\mathrm{bright}},|{-}\rangle^{\mathrm{bright}}\right\}.$ (64)
@@
$\begin{array}[]{rclclc}U_{\mathrm{E}}|{0}\rangle_{\mathrm{E}}|{0}\rangle_{\mathrm{A}}&=&\epsilon_{0,0}|{E_{0,0}}\rangle|{0}\rangle^{\mathrm{bright}}&+&\epsilon_{0,1}|{E_{0,1}}\rangle|{1}\rangle^{\mathrm{bright}}&+\\ &&\epsilon_{0,2}|{E_{0,2}}\rangle|{+}\rangle^{\mathrm{bright}}&+&\epsilon_{0,3}|{E_{0,3}}\rangle|{-}\rangle^{\mathrm{bright}},\\ \\ U_{\mathrm{E}}|{0}\rangle_{\mathrm{E}}|{1}\rangle_{\mathrm{A}}&=&\epsilon_{1,0}|{E_{1,0}}\rangle|{0}\rangle^{\mathrm{bright}}&+&\epsilon_{1,1}|{E_{1,1}}\rangle|{1}\rangle^{\mathrm{bright}}&+\\ &&\epsilon_{1,2}|{E_{1,2}}\rangle|{+}\rangle^{\mathrm{bright}}&+&\epsilon_{1,3}|{E_{1,3}}\rangle|{-}\rangle^{\mathrm{bright}}.\par\end{array}$ (65)
@@
$\begin{array}[]{c}\forall|{\psi}\rangle_{\mathrm{A}}=\sum_{i}\alpha_{i}|{i}\rangle_{\mathrm{A}},\quad\forall|{j}\rangle\in J_{{\operatorname{error}}}:\\[6.45831pt] \sum_{i,k}\alpha_{i}\epsilon_{i,k}\beta_{k,j}|{E_{i,k}}\rangle=0,\end{array}$ (66)
@@
$\sum_{i}\alpha_{i}\epsilon_{i,j}|{E_{i,j}}\rangle=0,$ (67)
@@
    $\varepsilon_{0,3}|{E_{0,3}}\rangle+\varepsilon_{1,3}|{E_{1,3}}\rangle=0.$ (70)
@@
    $\epsilon_{0,2}|{E_{0,2}}\rangle-\epsilon_{1,2}|{E_{1,2}}\rangle=0.$ (71)
@@
$\begin{array}[]{rcccccl}U_{\mathrm{E}}|{0}\rangle_{\mathrm{E}}|{0}\rangle_{\mathrm{A}}&=&\epsilon_{0,0}|{E_{0,0}}\rangle|{0}\rangle^{\mathrm{bright}}&+&\epsilon_{0,2}|{E_{0,2}}\rangle|{+}\rangle^{\mathrm{bright}}&+&\epsilon_{0,3}|{E_{0,3}}\rangle|{-}\rangle^{\mathrm{bright}},\\ U_{\mathrm{E}}|{0}\rangle_{\mathrm{E}}|{1}\rangle_{\mathrm{A}}&=&\epsilon_{1,1}|{E_{1,1}}\rangle|{1}\rangle^{\mathrm{bright}}&+&\epsilon_{1,2}|{E_{1,2}}\rangle|{+}\rangle^{\mathrm{bright}}&+&\epsilon_{1,3}|{E_{1,3}}\rangle|{-}\rangle^{\mathrm{bright}},\end{array}$ (72)
@@
Appendix G Examples of Attacks on QKD

### v3

@@ C.3. Attack Constraints
Appendix D Computation of Bright Illumination Attack Using Reversed-Space
In this appendix, we show the full computation of the Reversed-Space Attack discussed in Section 7.4, beginning from the behavior of the implementation that was discovered through Quantum Fuzzing in Section 7.3 and resulting in the original attack described in Section 7.1. Our modeling of the device will rely only on the information learned on the device from fuzzing, and on the fact that the device employs a single unitary transformation independently of the basis it wishes to measure.
@@ Appendix D Computation of Bright Illumination Attack Using Reversed-Space
The vulnerabilities shown in Section 7.3 show that Bob’s device has an unintended configuration that can be triggered by Eve (the linear mode of the APDs), and that in this configuration, his device suffers from a collection of Measurement Space Vulnerabilities and Interpretation Vulnerabilities (as defined in Section 3.2.1). We will now formally describe these vulnerabilities.
@@
Note that, according to the definition of high-photon-number states with diagonal polarization (see Appendix A), the overlap between “bright Hadamard” states and “bright computational” states is given, without loss of generality, by:
(58) ${\langle{+}|}^{\mathrm{bright}}|{0}\rangle^{\mathrm{bright}}=\left(\frac{1}{\sqrt{2^{k}}}\sum_{l=0}^{k}\sqrt{\binom{k}{l}}{\langle{l,k-l}|}_{\mathrm{H},\mathrm{V}}\right)|{0,k}\rangle_{\mathrm{H},\mathrm{V}}=\frac{1}{\sqrt{2^{k}}}=O\left(2^{-\frac{k}{2}}\right).$
@@
(59) $\mathcal{H}^{\mathrm{B}}=\operatorname{span}\left\{|{\psi_{0}}\rangle_{\mathrm{B}},|{\psi_{1}}\rangle_{\mathrm{B}},|{\psi_{2}}\rangle_{\mathrm{B}},|{\psi_{3}}\rangle_{\mathrm{B}}\right\}.$
@@
    (60) $J_{0}=\{|{\psi_{0}}\rangle_{\mathrm{B}}\},\quad J_{1}=\{|{\psi_{1}}\rangle_{\mathrm{B}}\}.$
@@
    (61) $J_{0}=\{|{\psi_{2}}\rangle_{\mathrm{B}}\},\quad J_{1}=\{|{\psi_{3}}\rangle_{\mathrm{B}}\}.$
@@
(62) $\mathcal{H}^{\mathrm{P}}=\operatorname{span}\left\{U_{\mathrm{B}}^{\dagger}|{\psi_{i}}\rangle_{\mathrm{B}}\;\big|\;i=0,1,2,3\right\}=\operatorname{span}\left\{|{0}\rangle^{\mathrm{bright}},|{1}\rangle^{\mathrm{bright}},|{+}\rangle^{\mathrm{bright}},|{-}\rangle^{\mathrm{bright}}\right\}.$
@@
(63) $\begin{array}[]{rclclc}U_{\mathrm{E}}|{0}\rangle_{\mathrm{E}}|{0}\rangle_{\mathrm{A}}&=&\epsilon_{0,0}|{E_{0,0}}\rangle|{0}\rangle^{\mathrm{bright}}&+&\epsilon_{0,1}|{E_{0,1}}\rangle|{1}\rangle^{\mathrm{bright}}&+\\ &&\epsilon_{0,2}|{E_{0,2}}\rangle|{+}\rangle^{\mathrm{bright}}&+&\epsilon_{0,3}|{E_{0,3}}\rangle|{-}\rangle^{\mathrm{bright}},\\ \\ U_{\mathrm{E}}|{0}\rangle_{\mathrm{E}}|{1}\rangle_{\mathrm{A}}&=&\epsilon_{1,0}|{E_{1,0}}\rangle|{0}\rangle^{\mathrm{bright}}&+&\epsilon_{1,1}|{E_{1,1}}\rangle|{1}\rangle^{\mathrm{bright}}&+\\ &&\epsilon_{1,2}|{E_{1,2}}\rangle|{+}\rangle^{\mathrm{bright}}&+&\epsilon_{1,3}|{E_{1,3}}\rangle|{-}\rangle^{\mathrm{bright}}.\par\end{array}$
@@
(64) $\begin{array}[]{c}\forall|{\psi}\rangle_{\mathrm{A}}=\sum_{i}\alpha_{i}|{i}\rangle_{\mathrm{A}},\quad\forall|{j}\rangle\in J_{{\operatorname{error}}}:\\[6.45831pt] \sum_{i,k}\alpha_{i}\epsilon_{i,k}\beta_{k,j}|{E_{i,k}}\rangle=0,\end{array}$
@@
(65) $\sum_{i}\alpha_{i}\epsilon_{i,j}|{E_{i,j}}\rangle=0,$
@@
    (68) $\varepsilon_{0,3}|{E_{0,3}}\rangle+\varepsilon_{1,3}|{E_{1,3}}\rangle=0.$
@@
    (69) $\epsilon_{0,2}|{E_{0,2}}\rangle-\epsilon_{1,2}|{E_{1,2}}\rangle=0.$
@@
(70) $\begin{array}[]{rcccccl}U_{\mathrm{E}}|{0}\rangle_{\mathrm{E}}|{0}\rangle_{\mathrm{A}}&=&\epsilon_{0,0}|{E_{0,0}}\rangle|{0}\rangle^{\mathrm{bright}}&+&\epsilon_{0,2}|{E_{0,2}}\rangle|{+}\rangle^{\mathrm{bright}}&+&\epsilon_{0,3}|{E_{0,3}}\rangle|{-}\rangle^{\mathrm{bright}},\\ U_{\mathrm{E}}|{0}\rangle_{\mathrm{E}}|{1}\rangle_{\mathrm{A}}&=&\epsilon_{1,1}|{E_{1,1}}\rangle|{1}\rangle^{\mathrm{bright}}&+&\epsilon_{1,2}|{E_{1,2}}\rangle|{+}\rangle^{\mathrm{bright}}&+&\epsilon_{1,3}|{E_{1,3}}\rangle|{-}\rangle^{\mathrm{bright}},\end{array}$
@@
Appendix E Examples of Attacks on QKD